OpenStudy (anonymous):

Can someone help me find the error? I'm supposed to "find the errors" in this assignment and I don't get how. Here's the first problem: The asymptotes are x= -3 and y= -1, thus h= -3 and k = -1 y= a/x+h + k y= 3/x+(-3) + (-1) y=3/x-3 -1

OpenStudy (ranga):

y = a / (x+h) + k If asymptote is at x = -3 that means -3 + h = 0 or h = 3 If another asymptote is at y = -1 then as x->infinity, y -> k and therefore k = -1 h = +3 and k = -1

OpenStudy (ranga):

y = a / (x + 3) - 1

OpenStudy (anonymous):

@ranga so the error is that h= +3 and not -3?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

@ranga i understand now. thank you!

OpenStudy (ranga):

Is the first line "a" or "3"? y= a/(x+h) + k or y = 3/(x+h) + k ?

OpenStudy (anonymous):

a

OpenStudy (ranga):

How did it become 3 in the second and third line?